# Rotatable and Movable Antenna-Enabled Near-Field Integrated Sensing and Communication

### v2

@@ Abstract
Integrated sensing and communication (ISAC) is regarded as a promising technology for next-generation communication networks. As the demand for communication performance significantly increases, extremely large-scale antenna arrays and tremendously high-frequency bands get widely applied in communication systems, leading to the expansion of the near-field region. Therefore, it is necessary to reevaluate the performance of ISAC in the near-field, especially considering the spherical wave near-field channel with effective aperture loss. On a parallel track, movable antennas (MAs) and six-dimensional MAs (6DMAs) are proposed as emerging technologies to improve the performance of communication and sensing. Based on such a background, this paper investigates the performance of ISAC systems in the near-field region, focusing on a novel system architecture that employs rotatable MAs (RMAs). The proposed RMA-enabled near-field ISAC system utilizes two rotatable antenna planes at the base station (BS) for both transmission and reception, allowing dynamic adjustment of antenna positions and rotations to enhance both communication and sensing capabilities. Additionally, a spherical wave near-field channel model with respect to RMAs’ rotations and positions is derived by considering the effective aperture loss. Two designs are explored: a sensing-centric design that minimizes the Cramér-Rao bound (CRB) with signal-to-interference-plus-noise ratio (SINR) constraints, and a communication-centric design that maximizes the sum-rate with a CRB constraint. To solve the formulated optimization problems, the paper proposes two alternating optimization (AO) based algorithms composed of the semidefinite relaxation (SDR) method and the particle swarm optimization (PSO) method. Numerical results demonstrate the convergence and effectiveness of the proposed algorithms and the superiority of the proposed setups for both sensing and communication performance compared to traditional fixed antenna systems, highlighting the potential of RMAs to enhance ISAC systems in near-field scenarios.
@@
Rotatable and movable antenna (RMA), integrated sensing and communications (ISAC), near-field, particle swarm optimization (PSO).
@@ I Introduction
The rapid advancement of communication technologies and the increasing demand for seamless connectivity have given rise to new paradigms in wireless networks. One such emerging concept is Integrated Sensing and Communication (ISAC), which converges the traditionally separate domains of communication and sensing into a unified framework [1]. ISAC stands out due to its ability to efficiently allocate time, frequency, power, and hardware resources for simultaneous communication and sensing tasks [2]. This capability is expected to optimize spectrum utilization and improve the performance of next-generation wireless systems. Furthermore, ISAC can be combined with other emerging technologies, such as reconfigurable intelligent surfaces (RIS) [3, 4] and fluid antenna [5, 6] to further enhance the performance of sensing and communication. As a result, ISAC presents significant research potential.
In light of recent advancements in wireless communication and sensing technologies, there is an increasing demand to meet the challenging requirements of next-generation wireless networks [7]. This typically necessitates the application of extremely large-scale antenna arrays and tremendously high- frequency bands. As a result, the Rayleigh distance $\frac{2D^{2}}{\lambda}$ with $D$ denoting the antenna aperture and $\lambda$ denoting the wavelength, is predicted to significantly increase, which means that the near-field region will correspondingly expand, encompassing distances of several hundred meters. When the user terminals (UTs) or the sensing targets are located in the near-field region, the propagation of electromagnetic waves requires a more accurate wave modeling approach, which differs from the conventional planar wave modeling adopted in far-field scenarios [8]. In addition to using spherical wave modeling, the effective antenna aperture loss should also be considered in the near-field channel model, otherwise, it may lead to unreasonable inaccuracies [7, 9]. Therefore, it is essential to reevaluate the performance of ISAC with appropriate channel modeling for near-field scenarios.
@@
In conventional multiple-input multiple-output (MIMO) systems, fixed-position antennas (FPAs) are widely configured and employed both for sensing and communication. Such fixed deployment of antennas limits the diversity and spatial multiplexing performance of MIMO systems because the channel variation in the spatial field is not fully utilized [10]. To overcome the inherent limitations of FPAs, movable antennas (MAs) and six-dimensional MAs (6DMAs) are introduced to wireless networks, which can fully exploit the spatial degrees of freedom (DoFs) of antenna arrays in communication or sensing scenarios [11, 12, 13, 14]. Compared to FPAs, transceivers equipped with movable antennas (MAs) can dynamically adjust the positions of antenna elements based on channel state information (CSI), thereby reshaping the channel conditions to boost communication performance, or reconfiguring the geometric properties of the antenna arrays to enhance sensing capability [15]. Building upon MAs, 6DMA leverages the DoFs in both the three-dimensional (3D) positioning and rotation of antenna planar. This enables adaptive antenna resource allocation based on long-term or statistical user distribution, ultimately boosting network capacity. Due to the advantageous properties of MA and 6DMA, there has recently been extensive research on various wireless sensing and ISAC systems aided by these technologies [15, 16, 17, 18, 19, 20]. The authors in [15] investigated a MA-aided ISAC system with RIS and maximized the minimum beampattern gain by jointly optimizing the transmit beamforming, the positions of MAs, and the phase shifts of RIS. In addition, the authors in [16] studied the CRB-minimization problem for a MA-aided ISAC system. Based on existing research on the use of movable antennas (MA) in ISAC systems considering far-field channel models [17, 18], the authors in [19] investigated the application of MAs in a dual-functional full-duplex ISAC system with near-field channel model. Moreover, the authors in [20] proposed a wireless sensing system aided by 6DMA and minimized the Cramér-Rao bound (CRB) for estimating the directions of arrival by jointly optimizing the 6DMAs’ positions and rotations at the base station (BS), comparing it with MAs for directive and isotropic antenna radiation patterns. However, it is worth noting that the performance of the MA planes’ rotations in near-field ISAC remains unexplored, especially considering that the existing literature has employed a spherical wave model, which fails to account for the loss in channel gain resulting from the effective antenna aperture [21, 22]. On the other hand, although 6DMA can achieve significant gains in both communication and sensing performance, its multiple rotatable surfaces require high hardware costs. Therefore, in this paper, we propose a rotatable MAs (RMAs) enabled near-field ISAC system, which can be regarded as a 6DMA-aided system with only two antenna surfaces and movable elements. The main contributions of this paper are summarized as follows:
@@
    We propose an RMA-enabled near-field ISAC system and employ a spherical wave channel model considering the effective antenna aperture loss. In the proposed scenario, the BS is equipped with two rotatable MA planes for transmission and reception respectively and all the MAs can move in large-size regions, serving multiple downlink UTs as well as performing sensing on one target.
@@
    We investigate the near-field ISAC performance by considering the sensing-centric design and the communication-centric design. For the sensing-centric design, we formulate a CRB minimization problem with signal-to-interference-plus-noise (SINR) constraints. And for the communication-centric design, we formulate a sum-rate maximization problem with SINR constraints and a CRB constraint.
@@
    Numerical results are present to demonstrate the convergence and effectiveness of the proposed algorithms. We also compare the performance of different setups for antennas at BS to study the individual effect and synergy of RMAs’ rotations and element position movement. Moreover, the simulation results of discrete rotations are shown to analyze the performance loss with a given number of quantization bits.
The rest of this paper is organized as follows. In Section II, our proposed conceptual framework of the RMA-enabled near-field ISAC system model is described by sequentially introducing the RMA model at BS with its constraints, the near-field channel model considering the effective aperture loss, the transmit ISAC signal model, and the adopted performance metrics. In Section III, a CRB minimization problem for sensing-centric design is formulated, and an AO-based algorithm is proposed to solve it. In Section IV, a sum-rate maximization problem for communication-centric design is characterized, and the corresponding AO-based algorithm is proposed. Numerical results are provided to verify the effectiveness of the proposed framework and algorithms in Section V, and a concise conclusion is provided in Section VI.
@@ II-A RMA-BS Model
In this paper, we consider an RMA-enabled near-field ISAC system, where a monostatic BS with RMAs communicates with $K$ single-FPA UTs and concurrently performs sensing on one target. As depicted in Fig. 1, we assume that the BS is equipped with $N_{\text{t}}$ transmit MAs and $N_{\text{r}}$ receive MAs. The transmit and receive MAs can move in two 2D planes, respectively, namely the transmit plane (TP) and the receive plane (RP). Without loss of generality, we assume that the MAs move within the $y$-$z$ plane of the local Cartesian coordinate system defined by the center of the corresponding plane, and the local positions of the $n_{\text{t}}$-th transmit MA and the $n_{\text{r}}$-th receive MA can be respectively represented as
@@
where $\mathcal{C}^{\text{t}}$ and $\mathcal{C}^{\text{r}}$ are the square moving regions of size $D\times D$ with the center points at the origin of corresponding local coordinates.
Figure 1: RMA aided near-field ISAC system.
@@
Given $\bm{\varpi}^{\text{t}}$ and $\bm{\varpi}^{\text{r}}$, the rotation matrix can be defined as [13]
$\mathbf{F}\left(\bm{\varpi}^{\text{p}}\right)=\begin{bmatrix}\cos{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}&\cos{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}&-\sin{\alpha^{\text{p}}}\\ F_{21}\left(\bm{\varpi}^{\text{p}}\right)&F_{22}\left(\bm{\varpi}^{\text{p}}\right)&\cos{\alpha^{\text{p}}}\sin{\beta^{\text{p}}}\\ F_{31}\left(\bm{\varpi}^{\text{p}}\right)&F_{32}\left(\bm{\varpi}^{\text{p}}\right)&\cos{\alpha^{\text{p}}}\cos{\beta^{\text{p}}}\end{bmatrix},$
where $F_{21}\left(\bm{\varpi}^{\text{p}}\right)=\sin{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}-\cos{\beta^{\text{p}}}\sin{\gamma^{\text{p}}}$, $F_{22}\left(\bm{\varpi}^{\text{p}}\right)=\sin{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}+\cos{\beta^{\text{p}}}\cos{\gamma^{\text{p}}}$, $F_{31}\left(\bm{\varpi}^{\text{p}}\right)=\cos{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}+\sin{\beta^{\text{p}}}\sin{\gamma^{\text{p}}}$, and $F_{32}\left(\bm{\varpi}^{\text{p}}\right)=\cos{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}-\sin{\beta^{\text{p}}}\cos{\gamma^{\text{p}}}$. Then the positions of the $n_{\text{t}}$-th transmit MA and the $n_{\text{r}}$-th receive MA in the global Cartesian coordinate system can be expressed as
@@
where $\mathbf{F}^{\text{p}}=\mathbf{F}\left(\bm{\varpi}^{\text{p}}\right)$, $\mathbf{C}^{\text{t}}$ and $\mathbf{C}^{\text{r}}$ denote the centers’ positions of TP and RP, respectively.
@@ II-A1 Rotation Constraints to Avoid Signal Reflection
As [13], the TP and RP must meet some rotation constraints to avoid mutual signal reflections between them, which can be expressed as
@@
with $\mathbf{u}^{\text{p}}=\mathbf{F}^{\text{p}}\left[1,0,0\right]^{\mathsf{T}}$ denoting the outward normal vector of the TP or RP, which ensures that all the antennas are positioned within the halfspace that consists of vectors forming obtuse angles with the normal vector of another plane.
@@ II-B Channel Model
Since the signals are observed by the receivers from distinct angles, the resultant effective antenna aperture varies w.r.t the positions of the antennas. The effective antenna aperture is determined by the product of the maximum value of the effective area and the projection of the array normal to the direction of the signal, with the resultant loss in channel power gain referred to as effective aperture loss [21]. Consequently, under free-space line-of-sight propagation, the channel power gain between the $n_{\text{p}}$-th antenna and the $k$-th UT/target is given by [9]
@@
denotes the free-space path loss and the effective aperture loss, respectively, with $\delta\left(\mathbf{p}_{k},\mathbf{s},\mathbf{u}^{\text{p}}\right)$ representing the sign function w.r.t. $\left(\mathbf{p}_{k}-\mathbf{s}\right)^{\mathsf{T}}\mathbf{u}^{\text{p}}$ for guaranteeing non-negativity of the projection. Given that the size of each antenna is significantly smaller than the propagation distance, i.e., $d_{k}\gg\sqrt{A},\forall k\in[0,K]$, the variation of the channel among different points $\mathbf{s}\in\mathcal{S}_{n_{\text{p}}}$ is negligible. Therefore, (11) can be approximately rewritten as
@@
where $\lambda=\frac{c}{f_{\text{c}}}$ is the signal wavelength with $f_{\text{c}}$ denoting the carrier frequency and $c$ denoting the speed of propagation.

### v3

@@ Abstract
The aim of this article is to investigate the performance of near-field integrated sensing and communication (ISAC) systems using rotatable movable antennas (RMAs). In the proposed RMA-enabled system, the positions and rotations of antennas at the base station (BS) are dynamically adjusted to enhance both communication and sensing capabilities. Two designs are explored: i) a sensing-centric design that minimizes the Cramér-Rao bound (CRB) with signal-to-interference-plus-noise ratio (SINR) constraints, and ii) a communication-centric design that maximizes the sum-rate with a CRB constraint. To solve the formulated optimization problems, two alternating optimization (AO)-based algorithms are proposed capitalizing on the semidefinite relaxation (SDR) method and the particle swarm optimization (PSO) method. Numerical results demonstrate that: i) the proposed RMA-enabled system outperforms the conventional fixed-position antenna and non-rotatable movable antenna systems in both sensing-centric and communication-centric designs and RMAs’ rotations show a higher performance gain in communication-centric design; ii) the proposed optimization methods achieve the Pareto boundary in both sensing-centric and communication-centric designs.
@@
Integrated sensing and communications (ISAC), near-field, particle swarm optimization (PSO), rotatable and movable antenna (RMA).
@@ I Introduction
Integrated Sensing and Communication (ISAC) is a promising technology in the evolution of the sixth generation (6G) wireless systems, which converges the traditionally separate domains of communication and sensing into a unified framework [1]. Due to its ability to efficiently allocate time, frequency, power, and hardware resources for simultaneous communication and sensing tasks [2], ISAC has been officially approved by IMT-2030 as one of six decisive usage scenarios for 6G networks and attracted increasing research attention. To further enhance the performance of sensing and communication, ISAC has been combined with multiple-antenna technologies, such as multiple-input multiple-output (MIMO) [3], to exploit additional spatial degrees of freedom (DoFs) and achieve greater array gains for both communication and sensing tasks. These improvements contribute to higher network throughput and improved sensing resolution. In general, the performance gains offered by antenna arrays scale approximately linearly with the aperture size. This insight has inspired the development of large-aperture array architectures, including reconfigurable intelligent surfaces (RIS) [4, 5], dynamic metasurfaces [6], holographic MIMO [7], and other related technologies, which have been widely adopted in ISAC systems.
However, the shift toward large-aperture arrays represents more than a mere quantitative increase in array size; it necessitates a paradigm shift from conventional far-field transmission to near-field transmission. Specifically, the expansion of antenna aperture enlarges the near-field region—an aspect often overlooked in traditional wireless systems. When user terminals (UTs) or sensing targets are located within the near-field region, electromagnetic wave propagation must be modeled using spherical wavefronts rather than the planar wave approximation typically used in far-field scenarios [8]. This transition calls for a fundamental reexamination of existing signal processing algorithms and system designs, thereby stimulating a growing body of research on near-field ISAC [9, 10].
I-A Prior Works
Benefiting from near-field propagation, near-field ISAC can obtain accurate resolution in spatial sensing, such as joint sensing of targets’ angle and distance information, while ensuring high communication performance. Nevertheless, most near-field ISAC systems are based on conventional fixed-position antennas (FPAs). Such deployment limits the diversity and spatial multiplexing performance of MIMO systems because the channel variation in the spatial field is not fully utilized [11].
@@ I-A Prior Works
To overcome the inherent limitations of FPAs, movable antennas (MAs) and six-dimensional MAs (6DMAs) are introduced to wireless networks, which can fully exploit the spatial DoFs of antenna arrays in communication or sensing scenarios [12, 13, 14, 15]. MA-enabled systems can dynamically adjust the positions of antennas, thereby reshaping the channel conditions to boost communication performance, or reconfiguring the geometric properties of the antenna arrays to enhance sensing capability [16]. Building upon MAs, 6DMA leverages the DoFs in both the three-dimensional (3D) positioning and rotation of antenna planar. Due to the advantageous properties of MA and 6DMA, there has recently been extensive research on various wireless sensing and ISAC systems enabled by these technologies [16, 17, 18, 19, 20, 21]. The authors in [16] investigated a MA-enabled ISAC system with RIS and maximized the minimum beampattern gain by jointly optimizing the transmit beamforming, the positions of MAs, and the phase shifts of RIS. In addition, the authors in [17] studied the Cramér-Rao bound (CRB) minimization problem for a MA-enabled ISAC system. Based on existing research on the MA-enabled far-field ISAC systems [18, 19], the authors in [20] investigated the application of MAs in a dual-functional full-duplex ISAC system with near-field channel model. Moreover, the authors in [21] proposed a wireless sensing system enabled by 6DMA and minimized the CRB for estimating the directions of arrival by jointly optimizing the 6DMAs’ positions and rotations at the base station (BS), comparing it with MAs for directive and isotropic antenna radiation patterns.
However, it is worth noting that the performance of the MA planes’ rotations in near-field ISAC remains unexplored, especially considering that the existing literature has employed a spherical wave model, which fails to account for the loss in channel gain resulting from the effective antenna aperture [22, 23]. On the other hand, although 6DMA can achieve significant gains in both communication and sensing performance, its multiple rotatable surfaces require high hardware costs.
I-B Contributions
Motivated by the above issues, we propose a rotatable MAs (RMAs) enabled near-field ISAC framework, which can be regarded as a 6DMA-enabled system with movable elements on only two planes. The main contributions of this paper are summarized as follows:
@@ I-B Contributions
    We propose an RMA-enabled near-field ISAC framework. In the proposed scenario, the BS is equipped with two rotatable MA planes for transmission and reception, respectively, and all the MAs can move in large-size regions, serving multiple downlink UTs as well as performing sensing on one target.
@@
    We investigate the near-field ISAC performance by considering the sensing-centric design and the communication-centric design. For the sensing-centric design, we formulate a CRB minimization problem with signal-to-interference-plus-noise (SINR) constraints. Regarding the communication-centric design, we formulate a sum-rate maximization problem with SINR constraints and a CRB constraint.
@@
    Numerical results are present to demonstrate the convergence and effectiveness of the proposed algorithms. We also compare the performance of different setups for antennas at BS to study the individual effect and synergy of RMAs’ rotations and element position movement. Moreover, we utilize the simulation results of discrete rotations to analyze the performance loss with a given number of quantization bits.
I-C Organization and Notations
The rest of this paper is organized as follows. In Section II, our proposed conceptual framework of the RMA-enabled near-field ISAC system model is described. In Section III, a CRB minimization problem for sensing-centric design is formulated, and an AO-based algorithm is proposed to solve it. In Section IV, a sum-rate maximization problem for communication-centric design is characterized, and the corresponding AO-based algorithm is proposed. Numerical results are provided to verify the effectiveness of the proposed framework in Section V, and a concise conclusion is provided in Section VI.
@@ II-A RMA-BS Model
We consider an RMA-enabled near-field ISAC framework, where a monostatic BS with RMAs communicates with $K$ single-FPA UTs while simultaneously sensing a nearby target, as depicted in Fig. 1. The BS is equipped with $N_{\text{t}}$ transmit MAs and $N_{\text{r}}$ receive MAs that can move in two 2D planes, namely the transmit plane (TP) and the receive plane (RP), respectively. Without loss of generality, the MAs are assumed to move within the $y$-$z$ plane of the local Cartesian coordinate system defined by the center of the corresponding plane. Therefore, the local positions of the $n_{\text{t}}$-th transmit MA and the $n_{\text{r}}$-th receive MA can be respectively represented as follows:
@@
where $\mathcal{C}^{\text{t}}$ and $\mathcal{C}^{\text{r}}$ denote the square moving regions of size $D\times D$ with the center points at the origin of corresponding local coordinates.
Figure 1: RMA-enabled near-field ISAC system.
@@
Given $\bm{\varpi}^{\text{t}}$ and $\bm{\varpi}^{\text{r}}$, the rotation matrix can be defined as follows [14]:
$\mathbf{F}\left(\bm{\varpi}^{\text{p}}\right)\triangleq\begin{bmatrix}\cos{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}&\cos{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}&-\sin{\alpha^{\text{p}}}\\ F_{21}\left(\bm{\varpi}^{\text{p}}\right)&F_{22}\left(\bm{\varpi}^{\text{p}}\right)&\cos{\alpha^{\text{p}}}\sin{\beta^{\text{p}}}\\ F_{31}\left(\bm{\varpi}^{\text{p}}\right)&F_{32}\left(\bm{\varpi}^{\text{p}}\right)&\cos{\alpha^{\text{p}}}\cos{\beta^{\text{p}}}\end{bmatrix},$
where
$F_{21}\left(\bm{\varpi}^{\text{p}}\right)=\sin{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}-\cos{\beta^{\text{p}}}\sin{\gamma^{\text{p}}},$ (5)
$F_{22}\left(\bm{\varpi}^{\text{p}}\right)=\sin{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}+\cos{\beta^{\text{p}}}\cos{\gamma^{\text{p}}},$ (6)
$F_{31}\left(\bm{\varpi}^{\text{p}}\right)=\cos{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\cos{\gamma^{\text{p}}}+\sin{\beta^{\text{p}}}\sin{\gamma^{\text{p}}},$ (7)
$F_{32}\left(\bm{\varpi}^{\text{p}}\right)=\cos{\beta^{\text{p}}}\sin{\alpha^{\text{p}}}\sin{\gamma^{\text{p}}}-\sin{\beta^{\text{p}}}\cos{\gamma^{\text{p}}}.$ (8)
As a result, the positions of the $n_{\text{t}}$-th transmit MA and the $n_{\text{r}}$-th receive MA in the global Cartesian coordinate system can be expressed as follows:
@@
where $\mathbf{F}^{\text{p}}=\mathbf{F}\left(\bm{\varpi}^{\text{p}}\right)$; $\mathbf{C}^{\text{t}}$ and $\mathbf{C}^{\text{r}}$ denote the centers’ positions of TP and RP, respectively.
@@ II-A1 Rotation Constraints to Avoid Signal Reflection
As per [14], the TP and RP must meet some rotation constraints to avoid mutual signal reflections between them, which can be expressed as follows:
@@
with $\mathbf{u}^{\text{p}}=\mathbf{F}^{\text{p}}\left[1,0,0\right]^{\mathsf{T}}$ denoting the outward normal vector of the TP or RP. This ensures that all the antennas are positioned within the halfspace that consists of vectors forming obtuse angles with the normal vector of another plane.
@@ II-B Channel Model
We assume that both the users and the target are located within the near-field region of the RMA, owing to its large aperture size. As discussed earlier, this necessitates the use of a spherical wavefront-based model to accurately characterize the wireless channel. Specifically, in the context of spherical-wave propagation, the signals received at different antennas arrive from distinct angles and distances. Consequently, the effective aperture of each antenna element varies depending on its spatial position relative to the source, which significantly impacts the array response and beamforming behavior. The effective antenna aperture is determined by the product of the maximum value of the effective area and the projection of the array normal to the direction of the signal, with the resultant loss in channel power gain referred to as effective aperture loss [22]. Therefore, under free-space line-of-sight propagation, the channel power gain between the $n_{\text{p}}$-th antenna and the $k$-th UT/target is given by [10]
@@
denote the free-space path loss and the effective aperture loss, respectively, with $\delta\left(\mathbf{p}_{k},\mathbf{s},\mathbf{u}^{\text{p}}\right)$ representing the sign function w.r.t. $\left(\mathbf{p}_{k}-\mathbf{s}\right)^{\mathsf{T}}\mathbf{u}^{\text{p}}$ for guaranteeing non-negativity of the projection. Given that the size of each antenna is significantly smaller than the propagation distance, i.e., $d_{k}\gg\sqrt{A},\forall k\in[0,K]$, the variation of the channel among different points $\mathbf{s}\in\mathcal{S}_{n_{\text{p}}}$ is negligible. Therefore, (15) can be approximated as
@@
where $\lambda=\frac{c}{f_{\text{c}}}$ is the signal wavelength with $f_{\text{c}}$ denoting the carrier frequency and $c$ denoting the speed of propagation. In contrast to conventional far-field channel, the phase in (20) is related to the distance between the transmit and receive antennas, rather than merely to the angle between them, and it is a nonlinear function of antennas’ positions, rather than a linear one.